OpenStudy (anonymous):

Which is a better buy? 4 ounce of toothpaste for $1.76 8 ounce of toothpaste for $2.80

OpenStudy (anonymous):

8 ounce

OpenStudy (jim766):

1.76/ 4 = ? 2.80 / 8 = ?

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

8 ounces is better to buy

OpenStudy (anonymous):

Because if you double the cost of 4 ounce you get $3.52

OpenStudy (anonymous):

That is because when you use the unit rate, 1.76/4 = 0.44

OpenStudy (anonymous):

Oh man, may have been wrong. Sorry for any inconvenience, but the answer may have been $2.80/8, as that equals 0.35. :(